OpenStudy (babybritt002):

Ethanol is 0.79 How many liter of ethanol contain 1.5 kg of ethanol? Explain how you go your answer

OpenStudy (cuanchi):

density of Ethanol is 0.79 g/mL = m/V 1.5 kg ethanol x (1000g /1kg) x ( 1mL / 0.79 g) x ( 1L /1000 ml) = XXX L